OpenStudy (wade123):

Find an equation for the nth term of the arithmetic sequence. a16 = 21, a17 = -1

OpenStudy (wade123):

an = 351 - 22(n + 1)

OpenStudy (wade123):

@ganeshie8

OpenStudy (anonymous):

c.d=a17-a16=-1-21=-22 a16=a+(16-1)d 21=a+15*-22 21=a-330 a=21+330=351 tn=a+(n-1)d =351+(n-1)*-22

OpenStudy (wade123):

so this ?351 - 22(n - 1)

OpenStudy (anonymous):

correct.

OpenStudy (wade123):

thanks a lot!

OpenStudy (anonymous):

yw